OpenStudy (anonymous):

Solve, using the substitution method. y – 2x = 8 16 + 4x = 2y (Points : 1) The solution is (24, 0). The solution is (1, 10). There is no solution. There are an infinite number of solutions.

OpenStudy (anonymous):

y - 2x = 8 --- we want to get y by itself, therefore, we want the -2x on the other side. To do this, we do the opposite. It is showing -2x, so we add 2x to both sides. y - 2x + 2x = 2x + 8 -- combine like terms y = 2x + 8 now that we have y = 2x + 8, we can sub 2x + 8 in for y in the other equation. 16 + 4x = 2y 16 + 4x = 2(2x + 8) -- distribute through the parenthesis 16 + 4x = 4x + 16 -- subtract 4x from both sides 16 + 4x - 4x = 4x - 4x + 16 -- combine like terms 16 = 16 if the solution ends like this, there are INFINITE SOLUTIONS if it would have ended like this : 16 = 14...then it would be no solutions. any questions ?

OpenStudy (anonymous):

i cant believe it but tht actually made complete sense u r a genius

OpenStudy (anonymous):

I don't know about that...it takes time and practice to learn all of this, but you seem smart, you will eventually catch on :)

OpenStudy (anonymous):

Solve, using the substitution method. y = 4x + 6 5x – y = 6 (Points : 1) There are an infinite number of solutions. There is no solution. (18, 12) (12, 54)

OpenStudy (anonymous):

may i get help with this 1

OpenStudy (anonymous):

of course....we already know that y = 4x + 6. So we sub in 4x + 6 for y in the other equation. Can you do that ?

OpenStudy (anonymous):

If they had the same slope they's be parallel and have no solution, so you can cross that one out too. Leaving just C and D as options.

OpenStudy (anonymous):

if the equations were the same they'd have infinitely many. But they're not so you can cross that out